FERMI Machine Status

Status

Run 33
Shift Users Shifts / LDM, TeraFERMI
Beam to Fel 1 - Padres
Next users dedicated injection
Last users dedicated injection 2018-02-01
Control Room Fermi (tel. 8800)
Info updated at

Linac / e-beam

RF plants ON
E-beam Energy 1.353 GeV
Bunch Charge 699 pC

Seed laser

Status OPEN
SLU Status laser ON
Pulse Energy 20.2 µJ
Wavelength 252.2 nm

FEL-1 I0 monitor charts (30 minutes)

FEL-1

Harmonic 4
Wavelength 63.05 nm
Polarization Linear Horizontal

FEL-1 I0 monitor

Intensity 114 µJ
RMS 6 %

Spectrometer

Wavelength ----
Bandwidth ----
Last Updated on Friday, 21 August 2015 15:46